# Enhancing Inverse Reinforcement Learning through Encoding Dynamic Information in Reward Shaping

### v2

@@ Abstract
Adversarial-based inverse reinforcement learning (IRL) has shown promising results using reward shaping under deterministic settings. However, it struggles in stochastic environments where existing theoretical results no longer apply, leading to degraded performance. To address this issue, we propose a novel maximum causal entropy based off-policy IRL method with transition-aware reward shaping framework. Our method integrates transition model estimation directly to learn stochastic-invariant rewards. We conduct a thorough theoretical analysis, establishing bounds on reward error and performance differences to validate the effectiveness of our method. The experimental results in continuous locomotion tasks (MuJoCo) show that our method can achieve superior performance in stochastic environments and competitive performance in deterministic environments, with significant improvement in sample efficiency, compared to existing baselines. Additionally, we extend our framework to high-dimensional vision-based tasks, where our method shows promising results on multiple stochastic Atari games. These results demonstrate that embedding transition awareness into reward learning is critical for robust IRL in realistic stochastic settings.
@@ 1 Introduction
The notable approaches within IRL are Generative Imitation Learning (GIL) methods that build upon maximum entropy framework (ziebart2008maximum). These methods frame imitation learning as a maximum likelihood estimation problem on trajectory distributions, converting the distribution into a Boltzmann distribution parameterized by rewards under deterministic environment settings (wu2024variational). This closely mirrors the distribution approximation found in generative models (finn2016connection; swamy2021moments). Thus, model-free deep IL approaches often follow generative model structures, such as GANs (ho2016generative; fu2017learning) or diffusion models (reuss2023goal; wu2024diffusing), and require extensive sampling for distribution/score function matching in on-policy fashion (orsini2021matters). Model-based IL frameworks have also emerged, where model-based framework is designed to provide estimation for gradient and planning, leading to innovative combinations such as gradient-based IRL with model predictive control (MPC) (das2021model) and end-to-end differentiable IRL frameworks for complex robotics tasks (baram2016model; baram2017end; sun2021adversarial; rafailov2021visual). However, these approaches primarily address deterministic settings and struggle when applied to stochastic environments.
@@
Existing IRL methods, rooted in their maximum entropy nature, mostly exclusively focus on learning "deterministic" reward techniques. These methods, however, face significant performance degeneration in stochastic environments, leading to risk-seeking behavior and increased data requirements (ziebart2010modeling). For example, an agent trained under the deterministic Markov Decision Process (MDP) might aim to imitate expert behavior by seeking high rewards, yet fail to account for the low probability of some transitions in stochastic MDP settings. This happens because, in stochastic environments, the assumption of maximum entropy no longer holds. The trajectory distributions are not aligned with a Boltzmann distribution solely parameterized by deterministic rewards. In this case, the dynamic information must also be incorporated into the formulation. There are two likely solutions. One is massive sampling to cover all possible outcomes, which is computationally expensive in large state action spaces (devlin2011theoretical; gupta2022unpacking). The other is changing from maximum entropy framework to maximum causal entropy framework, estimating the dynamics information, and integrating it into the reward design, making the reward "stochastic". Traditional reward design is usually based on state only $R(s_{t})$ (torabi2018generative), state-action pair $R(s_{t},a_{t})$ (blonde2019sample), or transition tuple $R(s_{t},a_{t},s_{t+1})$ (fu2017learning), where the information inputted can be thought as a deterministic sample piece under the stochastic setting. The challenge in stochastic environments calls for a different perspective of rewards – stochastic rewards absorbing the transition information. More detailed literature survey has been has been investigated in Appendix A.
@@
We begin by introducing the necessary preliminaries on Markov Decision Processes (MDPs) and inverse reinforcement learning (IRL). We then present our model-enhanced reward shaping method, along with its theoretical guarantee. Next, we describe the full model-enhanced IRL framework, including its derivation from the maximum causal entropy objective, and provide theoretical analysis on the reward error bound and performance difference bound. Experimental results are reported on a range of MuJoCo and Atari benchmarks, demonstrating the effectiveness of our approach. We conclude with a discussion of related work and a summary of our findings.
@@ 2 Preliminaries
$V^{soft}(s_{t})=log\sum_{a_{t}\in\mathcal{A}}\exp{Q^{soft}(s_{t},a_{t})},$ $\quad Q^{soft}(s_{t},a_{t})=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[V^{soft}(s_{t+1})|s_{t},a_{t}\right],$
@@ 3 Transition-Aware Reward Shaping
With the above definition, we can transfer the proof of policy invariant property of our designed reward shaping (Equation 1) to showing the equivalence of soft advantage functions, which is proved in the following theorem. The detailed proof can be found in Theorem B.1.
@@ 4.1 Adversarial Formulation of Reward Shaping
We bridge this adversarial formulation with the original MCE IRL problem. In the following proposition, we give a sketch of the proof to show the connection between the objective function of the discriminator and MCE IRL. Proof details can be found in C.1.
@@ 4.2 Algorithm Framework
In this section, we present the overall framework of Model-Enhanced IRL and illustrate how transition model training is incorporated into the learning loop. We use ensemble gaussian dynamic models to adopt the transition (yu2020mopo; yu2021combo; rigter2022rambo). The transition model is updated in each policy optimization iteration similar as model-based RL approaches (janner2019trust; hansen2022temporal; zhan2024state). At each iteration, the updated transition model is utilized for reward learning and synthetic data generation in eval mode, which is stored in the synthetic trajectory replay buffer. Unlike AIRL and GAIL, our framework operates in an off-policy fashion, where samples used for both discriminator and policy update are drawn from a combination of the environmental replay buffer and the synthetic replay buffer. An overview of our framework is shown in  Figure 1, and detailed algorithmic steps and parameters are provided in 1.
Sample Efficiency: To improve sample efficiency, we leverage the estimated transition model to generate $H$-steps synthetic trajectories data alongside real interaction data, facilitating policy optimization. Given that the estimated transition model is inaccurate at the beginning, we employ a dynamic ratio between real and synthetic data to prevent the model from being misled by unlikely synthetic transitions (janner2019trust; zhan2024state). Specifically, early-stage generated trajectories are not stored persistently, unlike real interactions which are fully stored in the off-policy environmental replay buffer. To maintain training stability, we use a synthetic replay buffer with a size that gradually increases as training progresses, ensuring a balanced inclusion of synthetic data over time. The growth rates of the data ratio and buffer size are adjusted based on the complexity of the transition model learning process and can be fine-tuned via hyper-parameters. Details can be found in implementation section of Appendix G. Distribution Shift: To mitigate distribution shift (lee2020addressing; lin2020model) during training, we employ a strategy involving the learned transition model. Typically, during interaction, the real state $s_{t}$ is used as input to the actor, and the resulting action $a_{t}$ is applied in the environment. To incorporate the transition model, we predict a synthetic state $\hat{s}_{t}$ from previous $s_{t-1}$ and $a_{t-1}$. This generated $\hat{s}_{t}$ is then fed into the actor to produce action $\hat{a}_{t}$. The actions $a_{t}$ and $\hat{a}_{t}$ are mixed and applied to the environment with a certain ratio, and the resulting pairs $(s_{t},a_{t})$ or $(s_{t},\hat{a}_{t})$ are stored in the environmental replay buffer. This approach helps balance the exploration of real and model-predicted dynamics, reducing the impact of distributional discrepancies.

### v3

@@ Abstract
Adversarial-based inverse reinforcement learning (IRL) has shown promising results using reward shaping under deterministic settings, but it struggles in stochastic environments. To address this issue, we propose a novel maximum causal entropy based off-policy IRL method with transition-aware reward shaping framework. Our method integrates transition model estimation directly to learn stochastic-invariant rewards. We conduct a thorough theoretical analysis, establishing bounds on reward error and performance differences to validate the effectiveness of our method. The experimental results in continuous locomotion tasks (MuJoCo) show that our method can achieve superior performance in stochastic environments and competitive performance in deterministic environments, with significant improvement in sample efficiency, compared to existing baselines. Additionally, we extend our framework to high-dimensional vision-based tasks, where our method shows promising results on multiple stochastic Atari games. These results demonstrate that embedding transition awareness into reward learning is critical for robust IRL in realistic stochastic settings.
@@ 1 Introduction
The notable approaches within IRL are Generative Imitation Learning (GIL) methods that build upon maximum entropy framework (ziebart2008maximum). These methods frame imitation learning as a maximum likelihood estimation problem on trajectory distributions, converting it into a Boltzmann distribution parameterized by rewards under deterministic environment settings (wu2024variational). This closely mirrors the distribution approximation found in generative models (finn2016connection; swamy2021moments). Thus, model-free deep IL approaches often follow generative model structures, such as GANs (ho2016generative; fu2017learning) or diffusion models (reuss2023goal; wu2024diffusing), and require extensive sampling for distribution/score function matching in on-policy fashion (orsini2021matters). Model-based IL frameworks have also emerged, where model-based framework is designed to provide estimation for gradient and planning, leading to innovative combinations such as gradient-based IRL with model predictive control (MPC) (das2021model) and end-to-end differentiable IRL frameworks for complex robotics tasks (baram2016model; baram2017end; sun2021adversarial; rafailov2021visual). However, these approaches primarily address deterministic settings and struggle when applied to stochastic environments.
@@
Existing IRL methods, rooted in their maximum entropy nature, mostly exclusively focus on learning "deterministic" reward techniques. These methods, however, face significant performance degeneration in stochastic environments, leading to risk-seeking behavior and increased data requirements (ziebart2010modeling). For example, an agent trained under the deterministic Markov Decision Process (MDP) might aim to imitate expert behavior by seeking high rewards, yet fail to account for the low probability of some transitions in stochastic MDP settings. This happens because, in stochastic environments, the assumption of maximum entropy no longer holds. The trajectory distributions are not aligned with a Boltzmann distribution solely parameterized by deterministic rewards. In this case, the dynamic information must also be incorporated into the formulation. There are two likely solutions: One is massive sampling to cover all possible outcomes, which is computationally expensive in large state action spaces (devlin2011theoretical; gupta2022unpacking); The other is changing from maximum entropy framework to maximum causal entropy framework, estimating the dynamics information, and integrating it into the reward design, making the reward "stochastic". Traditional reward design is usually based on state only $R(s_{t})$ (torabi2018generative), state-action pair $R(s_{t},a_{t})$ (blonde2019sample), or transition tuple $R(s_{t},a_{t},s_{t+1})$ (fu2017learning), where the information inputted can be thought as a deterministic sample piece under the stochastic setting. The challenge in stochastic environments calls for a different perspective of rewards – stochastic rewards absorbing the transition information. More detailed literature survey has been has been investigated in Appendix A.
@@
We begin by introducing the necessary preliminaries on Markov Decision Processes (MDPs) and inverse reinforcement learning (IRL). We then present our model-enhanced reward shaping method, along with its theoretical guarantee. Next, we describe the full model-enhanced IRL framework, including its derivation from the maximum causal entropy objective, and provide theoretical analysis on the reward error bound and performance difference bound. Experimental results are reported on a range of MuJoCo and Atari benchmarks, demonstrating the effectiveness of our approach. We conclude with a summary of our findings.
@@ 2 Preliminaries
$V^{soft}(s_{t})=\log\sum_{a_{t}\in\mathcal{A}}\exp{Q^{soft}(s_{t},a_{t})},$ $\quad Q^{soft}(s_{t},a_{t})=R(s_{t},a_{t})+\gamma\mathbb{E}_{\mathcal{T}}\left[V^{soft}(s_{t+1})|s_{t},a_{t}\right],$
@@ 3 Transition-Aware Reward Shaping
With the above definition, we can transfer the proof of policy invariant property of our designed reward shaping (Equation 1) to showing the equivalence of soft advantage functions, which is proved in the following theorem. The detailed proof can be found in Appendix B.
@@ 4.1 Adversarial Formulation of Reward Shaping
We bridge this adversarial formulation with the original MCE IRL problem. In the following proposition, we give a sketch of the proof to show the connection between the objective function of the discriminator and MCE IRL. Proof details can be found in Appendix C.
@@ 4.2 Algorithm Framework
In this section, we present the overall framework of Model-Enhanced IRL and illustrate how transition model training is incorporated into the learning loop. We use ensemble gaussian dynamic models to adopt the transition (yu2020mopo; yu2021combo; rigter2022rambo). The transition model is updated in each policy optimization iteration similar as model-based RL approaches (janner2019trust; hansen2022temporal; zhan2024state). At each iteration, the updated transition model is utilized for reward learning and synthetic data generation in eval mode, which is stored in the synthetic trajectory replay buffer. Unlike AIRL and GAIL, our framework operates in an off-policy fashion, where samples used for both discriminator and policy update are drawn from a combination of the environmental replay buffer and the synthetic replay buffer. An overview of our framework is shown in  Figure 1, and detailed algorithmic steps and parameters are provided in Algorithm 1.
Sample Efficiency:
To improve sample efficiency, we leverage the estimated transition model to generate $H$-steps synthetic trajectories data alongside real interaction data, facilitating policy optimization. Given that the estimated transition model is inaccurate at the beginning, we employ a dynamic ratio between real and synthetic data to prevent the model from being misled by unlikely synthetic transitions (janner2019trust; zhan2024state). Specifically, early-stage generated trajectories are not stored persistently, unlike real interactions which are fully stored in the off-policy environmental replay buffer. To maintain training stability, we use a synthetic replay buffer with a size that gradually increases as training progresses, ensuring a balanced inclusion of synthetic data over time. The growth rates of the data ratio and buffer size are adjusted based on the complexity of the transition model learning process and can be fine-tuned via hyper-parameters. Details can be found in Appendix G.
Distribution Shift:
To mitigate distribution shift (lee2020addressing; lin2020model) during training, we employ a strategy involving the learned transition model. Typically, during interaction, the real state $s_{t}$ is used as input to the actor, and the resulting action $a_{t}$ is applied in the environment. To incorporate the transition model, we predict a synthetic state $\hat{s}_{t}$ from previous $s_{t-1}$ and $a_{t-1}$. This generated $\hat{s}_{t}$ is then fed into the actor to produce action $\hat{a}_{t}$. The actions $a_{t}$ and $\hat{a}_{t}$ are mixed and applied to the environment with a certain ratio, and the resulting pairs $(s_{t},a_{t})$ or $(s_{t},\hat{a}_{t})$ are stored in the environmental replay buffer. This approach helps balance the exploration of real and model-predicted dynamics, reducing the impact of distributional discrepancies.